### v2

# GeoCAD: Local Geometry-Controllable CAD Generation with Large Language Models

Abstract
Local geometry-controllable computer-aided design (CAD) generation aims to modify local parts of CAD models automatically, enhancing design efficiency. It also ensures that the shapes of newly generated local parts follow user-specific geometric instructions (e.g., an isosceles right triangle or a rectangle with one corner cut off). However, existing methods encounter challenges in achieving this goal. Specifically, they either lack the ability to follow textual instructions or are unable to focus on the local parts. To address this limitation, we introduce GeoCAD, a user-friendly and local geometry-controllable CAD generation method. Specifically, we first propose a complementary captioning strategy to generate geometric instructions for local parts. This strategy involves vertex-based and VLLM-based captioning for systematically annotating simple and complex parts, respectively. In this way, we caption $\sim$221k different local parts in total. In the training stage, given a CAD model, we randomly mask a local part. Then, using its geometric instruction and the remaining parts as input, we prompt large language models (LLMs) to predict the masked part. During inference, users can specify any local part for modification while adhering to a variety of predefined geometric instructions. Extensive experiments demonstrate the effectiveness of GeoCAD in generation quality, validity and text-to-CAD consistency. Code will be available at https://github.com/Zhanwei-Z/GeoCAD.
Figure 1: Local geometry-controllable CAD generation achieved by GeoCAD. The input comprises: (1) an original CAD model (the left side), (2) the local part to be modified (highlighted in blue), and (3) user-specific geometric instructions. Subsequently, GeoCAD outputs the revised CAD models where only the target part is altered while adhering to the provided geometric instructions.
1 Introduction
Computer-Aided Design (CAD) is pivotal in industrial design, driving innovation and efficiency across diverse domains such as mechanical manufacturing cherng1998feature ; ganin2021computer ; khan2024cad . In CAD tools (such as SolidWorks and AutoCAD), the sketch-extrude-modeling (SEM) workflow shahin2008feature ; wu2021deepcad ; pmlr-v162-xu22k ; pmlr-v202-xu23f is commonly employed, enabling users to control the parametric design process effectively. During this process, users sequentially extrude each 2D sketch into 3D shapes to construct complex solid CAD models, with each sketch comprising one or multiple local loops. Each local loop typically represents a pattern or geometric shape, serving as the fundamental closed-path element of a CAD model pmlr-v162-xu22k ; pmlr-v202-xu23f .
In practice, any minor mistake in local parts (i.e., local loops [^1] In the following, local parts refer to local loops, the finest-grained closed-path elements of a sketch.) of a CAD model can potentially result in significant systemic errors. Thus, after drawing a draft CAD model, users generally need to modify its local parts to ensure that the final CAD product meets the expected functional or aesthetic requirements. Compared to manual modifications, if a deep-learning method can automatically adjust the shapes of local parts according to user-defined geometric instructions  [^2] In this paper, geometric instructions denote textual captions of loop shapes. (e.g., an isosceles right triangle or a rectangular shape with one corner removed), it would significantly reduce labor costs in CAD product optimization. Moreover, such a method must retain the remaining CAD parts unchanged while ensuring that the newly generated local parts integrate with them without conflict. We refer to these capabilities as local geometry-controllable CAD generation.
Unfortunately, existing controllable CAD generation methods face challenges in achieving local geometry-control. Specifically, pmlr-v162-xu22k ; pmlr-v202-xu23f ; wu2024cadvlm ; li2025revisiting ; zhang2025diffusion typically take partial CAD parts or attributes (e.g., incomplete sketches, topological or geometric parameters) as input and automatically generate new CAD models. Yet, they lack the ability to follow textual instructions, which hinders users from expressing their requirements intuitively and conveniently. To resolve this limitation, some text-to-CAD methods based on LLMs or transformers vaswani2017attention have demonstrated meaningful progress li2024cad ; khan2024text2cad ; xu2024cad ; yuan2025cad ; wang2025cad ; wang2025text ; alrashedy2025generating ; zhang2025flexcad . However, these methods are not applicable for local geometry-controllable generation. Specifically, li2024cad ; khan2024text2cad ; xu2024cad ; wang2025cad ; wang2025text ; alrashedy2025generating typically generate a new CAD model from scratch based on textual instructions, making it difficult to fully focus on the required local parts. In addition, khan2024text2cad ; yuan2025cad ; wang2025text ; wang2025cad primarily collect textual descriptions of CAD models from global 3D views rather than local 2D views. These 3D views are generally oblique, which prevents capturing accurate geometric attributes (such as length and angle) of local parts for training.  zhang2025flexcad can focus on local parts well but incorporates little geometric constraint, thereby struggling to follow geometric instructions.
In this paper, we propose GeoCAD, a user-friendly and local geometry-controllable CAD generation method. As shown in Fig. 1, GeoCAD takes the original CAD model, the local parts (highlighted in blue), and user-specific geometric instructions as inputs. The local parts are then generated by GeoCAD to align with the instructions, and are combined with the remaining parts to create new CAD models. To achieve this objective, the primary challenge is addressing the insufficiency of training data, specifically the geometric instructions for local parts. Given that manual captioning is prohibitively costly and labor-intensive, we introduce a complementary captioning strategy. Specifically, we categorize local parts into simple and complex groups based on their internal side types and numbers. Simple parts correspond to common geometric shapes (e.g., triangles with three lines, quadrilaterals with four lines), while complex parts typically represent more intricate visual patterns. For complex parts, we render them as 2D images and then employ advanced vision large language models (VLLMs) achiam2023gpt ; Qwen-VL to derive descriptive captions. However, for simple parts, VLLMs fail to achieve accurate fine-grained captioning. For example, VLLMs do not reliably distinguish whether a quadrilateral is a rhombus based solely on an image. To overcome this limitation, we introduce a vertex-based captioning method for simple parts. This involves extracting vertex coordinates from the original CAD model and then analyzing geometric attributes for accurate classification. For instance, if a quadrilateral has four lines of equal length, it is a rhombus; if it contains right angles, it is further categorized as a square. Utilizing the complementary strategy, we have successfully captioned approximately 221k different local parts, comprising 116k complex parts and 105k simple parts. Inspired by the success of LLMs in text-to-CAD generation xu2024cad ; alrashedy2025generating ; yuan2025cad ; wang2025cad ; wang2025text ; zhang2025flexcad , during training, given a CAD model, we randomly mask a local part and prompt LLMs to predict this part using the corresponding geometric instruction and the remaining visible parts as inputs. Once trained, in real-world applications, users can mask any local part for modification based on various geometric instructions. The new local parts generated by GeoCAD are then integrated with the remaining parts of the original CAD model to form new CAD models. Overall, our contributions are:
  • We propose GeoCAD, a local geometry-controllable CAD generation method, enabling users to express design intent for specific parts through geometric instructions.
  • To the best of our knowledge, GeoCAD is the first to achieve local geometry-control in the CAD generation field. To achieve this, we propose a complementary captioning pipeline to annotate $\sim$221k distinct local parts for the following two-stage LLM fine-tuning.
  • Extensive experiments demonstrate that GeoCAD significantly enhances generation quality, validity, and text-to-CAD consistency in local geometry-controllable CAD generation.
2 Related Work
CAD Model Generation. Existing CAD generation methods can be categorized into three types: constructive solid geometry (CSG), boundary representation (B-rep) and sketch-and-extrude modeling (SEM). CSG constructs CAD models by combining primitives (e.g., cubes or spheres) into a tree laidlaw1986constructive ; chen2024fr ; yu2024d ; romano2025point . B-rep denotes CAD models as interconnected faces, edges, and vertices ansaldi1985geometric ; cherenkova2024spelsnet ; xu2024brepgen ; shen2025mesh2brep . Compared to CSG and B-rep, SEM-based methods wu2021deepcad ; pmlr-v162-xu22k ; pmlr-v202-xu23f ; khan2024cad ; ma2024draw ; yuan2025cad ; wang2025text ; wang2025cad ; li2025revisiting ; zhang2025flexcad ; zhang2025diffusion ; chen2025cadcrafter are consistent with prevailing CAD tools, allowing users to sequentially extrude sketches into 3D shapes, with each sketch comprising one or multiple loops. Notably, within a sketch, any loop nested inside another loop serves as a hole. Recently, SEM-based controllable CAD generation has garnered a lot of attention due to its potential to revolutionize the design process pmlr-v162-xu22k ; pmlr-v202-xu23f ; wu2024cadvlm ; li2025revisiting ; zhang2025diffusion . Specifically, these methods allow for some level of control over the parts or attributes of the original CAD models. Among them, pmlr-v162-xu22k ; wu2024cadvlm ; li2025revisiting ; zhang2025diffusion achieve sketch-level control, while pmlr-v202-xu23f offers finer-grained control over local loops. Despite these capabilities, these methods struggle to follow textual instructions, limiting users from conveying their design intent in an intuitive and convenient manner.
On the other hand, current text-to-CAD methods that have demonstrated meaningful progress li2024cad ; khan2024text2cad ; xu2024cad ; yuan2025cad ; wang2025cad ; wang2025text ; alrashedy2025generating . Notably, li2024cad ; khan2024text2cad ; xu2024cad ; wang2025cad ; wang2025text ; alrashedy2025generating typically generate a new CAD model from the ground up based on textual instructions, which limits their ability to precisely target or refine specific local parts as per user specifications. Moreover, khan2024text2cad ; yuan2025cad ; wang2025text primarily gather textual descriptions from global 3D perspectives rather than localized 2D views. These 3D perspectives are typically captured in oblique orientations, which limits their ability to precisely quantify critical geometric attributes (e.g., length and angle) of local parts during the training process.  zhang2025flexcad can effectively concentrate on the generation of local parts but fails to follow geometric instructions. To sum up, current studies lack the ability to achieve local geometry-controllable generation.
Large Language Models (LLMs). Compared to traditional deep-learning based models zhou2024learning ; zhang2025straj ; dai2024harmonious ; dong2024global ; shen2024imagpose ; li2025difiisr , LLMs have recently demonstrated a remarkable ability to follow textual instructions touvron2023llama ; achiam2023gpt ; yuan2024instance ; yang2024qwen2 ; liuyue_efficient_reasoning ; das2025security ; liuyue_GuardReasoner-VL . Leveraging this capability, LLMs have shown notable versatility and efficacy across diverse applications zhang2024llamaadapter ; li2024finetuning ; yang2025lidar ; fan2025improving ; zhang2025collm ; shojaee2025llmsr . Users can employ various textual instructions to direct LLMs in accomplishing diverse tasks like code generation gu2025effectiveness ; dong2025codescore and question answering singhal2025toward ; kuang2025natural . As a branch, vision large language models (VLLMs) have also achieved significant success in vision domains liu2024visual ; zang2025contextual ; li2025uni . More recently, both LLMs and VLLMs have shown promise in CAD generation xu2024cad ; alrashedy2025generating ; yuan2025cad ; wang2025cad ; wang2025text ; zhang2025flexcad . Specifically, xu2024cad ; yuan2025cad ; wang2025text ; wang2025cad ; zhang2025flexcad primarily rely on VLLMs for CAD caption synthesis or fine-tune LLMs or transformers vaswani2017attention to generate CAD models from scratch. On the other hand, alrashedy2025generating employs a training-free manner to generate CAD codes via informative prompts. As mentioned above, these methods either cannot effectively focus on local generation or struggle to follow geometric instructions accurately. Distinguished from them, our GeoCAD excels in local part generation while precisely adhering to geometric instructions.
3 Methodology
In this section, we present GeoCAD, a user-friendly and local geometry-controllable CAD generation method. As shown in Fig. 1, GeoCAD incorporates three inputs: (1) an original CAD model, represented in a hierarchically textual format proposed by FlexCAD zhang2025flexcad , (2) the local part designated for modification, and (3) geometric instructions specified by the user. GeoCAD then generates new CAD models, altering only the designated local part while closely adhering to the provided instructions. To achieve this, we first propose a complementary captioning strategy to generate $\sim$221k geometric instructions for local parts (Sec. 3.1). Building on these instructions, we then formulate a two-stage training pipeline to fine-tune LLMs for local CAD generation (Sec. 3.2).
3.1 Complementary Captioning for Local Parts
Figure 2: The complementary captioning strategy. (a) Vertex-based captioning for simple local parts. Vertex coordinates are initially extracted, followed by geometric analysis to enable precise captions. (b) VLLM-based captioning for complex local parts. We first convert complex parts into 2D images and subsequently employ powerful VLLMs to produce descriptive captions.
The main challenge in achieving local geometry-control is tackling the lack of training data, particularly concerning geometric instructions for local parts within 3D CAD models. Since manual captioning is excessively expensive and labor-intensive, we propose a complementary captioning strategy. In the beginning, we collect local parts (i.e., local loops) from the CAD models within the DeepCAD dataset wu2021deepcad , filtering out duplicates and discarding invalid ones (i.e., those that are not closed loops or involve intersecting line segments). Subsequently, we adopt the textual format introduced in FlexCAD zhang2025flexcad to represent CAD models and their local parts, where each local part is denoted as a contiguous string comprising the side type and vertex coordinates, as illustrated in Fig. 2. These local parts are then categorized into simple and complex groups based on their internal side numbers and types. Specifically, as shown in the lower part of Fig. 2(a), simple parts represent common geometric shapes (e.g., triangles with three lines, quadrilaterals with four lines, sectors with two lines and an arc), making up roughly 50% of the entire set of local parts, while complex parts typically exhibit more intricate visual patterns as shown in the lower part of Fig. 2(b).
As shown in the upper part of Fig. 2(b), for complex parts, we transform them into 2D images and then leverage powerful VLLMs achiam2023gpt ; Qwen-VL to obtain their geometric instructions (see the detailed prompts to guide VLLMs in the appendix). However, VLLMs exhibit limitations in fine-grained geometric descriptions for simple parts. For instance, they struggle to reliably discern whether a quadrilateral is a rhombus according to an image alone. To address this problem, we propose a vertex-based captioning method for simple parts. As shown in the upper part of Fig. 2(a), we first extract vertex coordinates from the original CAD text representation and then analyze geometric properties to precisely categorize these parts. For instance, given a quadrilateral, we can calculate its side lengths and inter-side angles based on its vertex coordinates. If it has four lines of equal length, it is a rhombus; if it includes right angles, it is further categorized as a square. Moreover, for partial simple parts, we also incorporate key dimensional parameters into the captions (such as the radius length of a circle and the side length of a square). In total, we annotate nearly 221k distinct local parts, consisting of 116k complex parts and 105k simple parts.
3.2 Fine-tuning LLMs with Geometric Instruction
With the geometric instructions derived in Sec. 3.1, we fine-tune LLMs to achieve local geometry-controllable CAD generation. The training procedure comprises two stages:
Stage 1: Pre-training for CAD-text Alignment (Optional). As mentioned in Sec. 3.1, we follow zhang2025flexcad to represent local parts using their internal side types and vertex coordinates. Since such CAD-specific geometric representation is typically absent from the pretraining corpus of LLMs, this stage focuses on aligning the representation of local parts with textual geometric instructions, thereby further enhancing the LLMs’ understanding of the CAD-specific representation. Specifically, as illustrated in Fig. 3, for each local part, we apply random data augmentation via translation, scaling, rotation, and reflection. Notably, the geometric instructions of augmented samples remain unchanged due to the geometric consistency (e.g., the geometric instructions of the augmented samples in Fig. 3 are all right trapezoids). Subsequently, for the initial and augmented samples, their corresponding instructions and answers are all employed to fine-tune LLMs.
Figure 3: The prompt template used in stage 1. Local parts are first augmented through translation, scaling, rotation, and reflection. Subsequently, we construct the corresponding prompt that incorporates the geometric instruction, and ask LLMs to predict both the initial and augmented parts.
Figure 4: The prompt template used in stage 2. Given a local part (highlighted in blue) in a CAD model, we formulate the prompt that integrates the geometric instruction (highlighted in green) and the remaining parts of the CAD model, and require LLMs to predict this local part.
Figure 5: The overall framework of GeoCAD. (a) Training process. Given a CAD model, we randomly mask a local loop within it. During stages 1 and 2, we design the corresponding prompts (as introduced in Fig. 3 and Fig. 4), and fine-tune LLMs. (b) Inference process. Users can optionally mask any local part for modification, driven by various geometric instructions (GI). The mask part is then infilled with the predicted local parts to construct new CAD models.
Stage 2: Instruction Fine-Tuning for Local Geometry-Control. In practice, when modifying a specific part of a CAD model, it is crucial to retain the other parts of the CAD model unchanged. Additionally, the newly generated part should integrate with them without any conflicts. To this end, inspired by FlexCAD zhang2025flexcad , at each epoch, for a given CAD model, we randomly mask a local part and design geometric instructions. These instructions are employed to prompt LLMs to predict this masked part autoregressively. However, FlexCAD’s training process has one critical limitation: its prompts lack geometric constraints during training. Consequently, once trained, FlexCAD struggles to follow geometric instructions. In light of this, as shown in Fig. 4, our prompts incorporate the geometric instructions as constraints when fine-tuning LLMs to generate predictions. As shown in Fig. 5, during stages 1 and 2, the cross-entropy (CE) loss between the predicted tokens and the answer tokens is back-propagated to update the trainable parameters of LLMs. Furthermore, we follow FelxCAD zhang2025flexcad to fine-tune the LLM using LoRA hu2022lora , which enables partial parameters training while freezing most parameter weights. This strategy allows us to retain the advantages of large-scale pre-trained models while accelerating convergence during optimization.
Inference. In practical applications, users can selectively mask any local part for modification, guided by various geometric instructions (e.g., a square, a rounded rectangle, or an annular sector). The mask part is then replaced with the predicted local parts, which are seamlessly integrated with the remaining parts of the original CAD model to form new CAD models, as shown in Fig. 5(b).
4 Experiments
4.1 Experimental Setup
Datasets. To maintain consistency with prior research zhang2025flexcad , we evaluate our GeoCAD on DeepCAD wu2021deepcad , a large-scale 3D sketch-extrude-modeling CAD dataset. This dataset contains 178,238 sketch-extrusion sequences, which are randomly partitioned into training, validation, and test subsets at a 90%-5%-5% ratio. Following established preprocessing protocols from SkexGen pmlr-v162-xu22k , we first eliminate duplicate and invalid sequences to ensure data quality. Subsequently, we follow FlexCAD zhang2025flexcad to convert the remaining CAD sequences into concise textual representations, which can be easily fed into LLMs. Within this dataset, we systematically collect and caption approximately 221k distinct local parts, including 116k complex parts and 105k simple parts.
Implementation Details. To ensure a fair comparison with FlexCAD zhang2025flexcad , we adopt Llama-3-8B meta2024introducing as the base LLM, which achieves competitive performance among open-source LLMs. We use the same LoRA hu2022lora setting as used in zhang2025flexcad , with a rank of 8 and an alpha of 32. In stage 1, we implement translation, scaling, rotation, and reflection for simple parts, while applying only translation and scaling to complex parts to avoid semantic inconsistencies in captions. The model is trained on 8 A100 GPUs using AdamW loshchilov2018decoupled , with a batch size of 32, a cosine annealing learning rate initialized at $5\times 10^{-4}$, and trained for 10 and 30 epochs across stage 1 and stage 2. During inference, we set the temperature $\tau$ and $\mathrm{Top\text{-}p}$ at 0.9 and 0.9 to balance quality and validity in local generation.
Metrics. As this work pioneers local geometry-controllable CAD generation, we propose a comprehensive evaluation benchmark based on three key aspects:
1) Generation quality. We adopt metrics from prior work pmlr-v162-xu22k ; pmlr-v202-xu23f ; zhang2025flexcad . Specifically, Coverage (COV) measures the diversity of generated shapes and helps identify whether the model suffers from mode collapse. Minimum Matching Distance (MMD) reports the average minimum distance between real data and the generated set. Jensen-Shannon Divergence (JSD) quantifies the similarity between the distributions of real and generated samples. Together, these metrics measure generation quality on generated CAD models with respect to the test set.
2) Validity. Predicted local parts must form closed loops and must not contain intersecting line segments. In addition, these parts should seamlessly integrate with the existing parts to enable successful rendering into valid 3D shapes, rather than invalid or empty outputs. Following zhang2025flexcad , we use Prediction Validity (PV) to quantify the overall validity rate of the generated predictions.
3) Text-to-CAD consistency. The generated 2D local parts should be consistent with user-defined geometric instructions. To measure this, we propose a vertex-based score (Ver-score) for assessing simple parts, and a VLLM-based score (VLLM-score) to evaluate complex parts. Finally, Realism denotes the human evaluation score, manually assessing whether the generated 3D CAD models fully satisfy user requirements for local geometry-control. See details of these metrics in the appendix.
4.2 Performance Comparision with Existing Methods
Baselines. As discussed above, most controllable CAD generation methods are not applicable to the local geometry-control task. Thus, we compare our GeoCAD with OpenAI-o3 o3 , one of the most powerful closed-source LLMs, and FlexCAD zhang2025flexcad , a state-of-the-art baseline for local CAD generation by fine-tuning LLMs. Without fine-tuning, the output format of the vanilla OpenAI-o3 model does not conform to the textual representation defined in zhang2025flexcad , making it unable to directly generate local parts. To address this, we improve the performance of OpenAI-o3 with a few-shot learning strategy. Moreover, we manually enhance FlexCAD’s performance when generating simple parts by adjusting the internal curve types and numbers. For example, when aiming to generate an isosceles trapezoid, we try our best to guide FlexCAD to produce a loop composed of four lines.
Table 1: Performance comparison on the DeepCAD test set. Five-shot denotes that each prompt includes five exemplars selected from the training set that are either identical or semantically similar to the target instruction. Exemplars used in OpenAI-o3 consist of instructions and answers, following the format shown in Fig. 4. Best performances are in bold, and the second-bests are marked by *.
[…]
OpenAI-o3 (five-shot) 53.6% 1.64 1.49 65.7% 33.6% 22.1% 18.7%
FlexCAD 58.3% 1.40 1.58 86.7% 19.8% 6.93% 13.6%
FlexCAD (five-shot) 59.4% 1.37 1.34 88.1% 43.5% 26.8% 20.2%
GeoCAD 64.9%* 1.13 0.98* 90.5%* 76.4%* 65.7%* 40.9%*
GeoCAD (five-shot) 66.0% *1.16 0.80 92.3% 82.2% 68.2% 43.6%
Figure 6: Qualitative comparison results for three methods. On the left, we show the original CAD models (in textual format), with the local parts to be modified (highlighted in blue, the same below). On the right, the upper section presents the user-defined geometric instructions, and the lower section displays the corresponding newly generated CAD models. Both FlexCAD and OpenAI-o3 are enhanced using five-shot learning. Red boxes indicate frequently occurring shapes in the training set (e.g., circles or rectangles) that do not conform to the given geometric instructions. Green boxes highlight local parts that are poorly integrated with the remaining parts of the original CAD models.
Quantitative Results. We randomly sampled 1k CAD models from the test set. For each CAD model, a local part was randomly masked, and each method was prompted to generate 10 new parts using 5 simple and 5 complex geometric instructions. Here, simple and complex instructions correspond to the generation of simple and complex local parts, respectively. After infilling, this process yielded a total of 10k generated CAD models per method. To compute the COV, MMD, and JSD metrics, which rely on a subset of ground-truth samples, we randomly selected 3k CAD models from the test set and calculated the average results over three separate runs. As presented in Table 1, OpenAI-o3 delivers subpar performance without fine-tuning, even when supported by five-shot learning. In comparison, our proposed GeoCAD achieves superior results over the state-of-the-art baseline, FlexCAD, particularly in terms of Ver-score, VLLM-score, and Realism, with significant improvements of up to 38.7%, 41.4%, and 23.4%, respectively. This is mainly because FlexCAD lacks the ability to align with geometric instructions during the generation of local parts. On the other hand, the few-shot learning ability of LLMs leads to performance improvements for both FlexCAD and our GeoCAD. Overall, the results demonstrate the clear advantage of our GeoCAD in generation quality, validity, and text-to-CAD consistency.
Figure 7: Additional qualitative results for GeoCAD. On the right, the upper section shows the user-defined instructions, while the lower section presents the newly generated CAD models.
Figure 8: GeoCAD is capable of precisely controlling the key dimensional parameters. The right side displays the newly generated CAD models and the corresponding geometric instructions.
Figure 9: Generalization ability of GeoCAD. On the right, each row contains two groups, with each group comprising three examples generated based on semantically similar instructions.
Qualitative Results. To intuitively demonstrate performance, we randomly selected six CAD models from the test set. As shown in Fig. 6, the results clearly highlight that our GeoCAD significantly improves controllability and text-to-CAD consistency compared to existing baseline methods. In particular, GeoCAD is able to modify local parts in a way that closely adheres to user-defined geometric instructions. In contrast, FlexCAD struggles to comply with such instructions and frequently generates overly common shapes, such as circles or rectangles (see green boxes in Fig. 6). Moreover, as shown in the red boxes in Fig. 6, both OpenAI-o3 and FlexCAD often produce local parts that fail to align properly with the remaining parts of the original CAD models, resulting in outputs that are functionally or aesthetically implausible. These visualizations further validate the superior local controllability and effectiveness of our proposed GeoCAD.
Furthermore, we provide additional qualitative results generated by GeoCAD. As illustrated in Fig. 7, given a CAD model, GeoCAD is capable of effectively modifying any target loop within it to form simple or complex patterns. Moreover, for certain simple parts, we incorporate specific dimensional constraints into the instructions, such as the radius of a circle, the side length of a square, and the length and width of a rectangle. As shown in Fig. 8, GeoCAD not only accurately generates the desired shapes but also adheres closely to the specified dimensional parameters. On the other hand, as shown in Fig. 9, GeoCAD demonstrates robust generalization capabilities in accurately understanding and executing semantically similar instructions, even when some of these instructions (e.g., a narrow, rounded-end rectangle and a right-angled triangle) never appeared in the training data.
4.3 Ablation Studies
We conduct a series of ablation studies under the same experimental settings described in Table 2.
Table 2: Effectiveness analysis of the complementary captioning strategy and pre-training. $\mathrm{Only\ Vertex\text{-}based\ Captioning}$ and $\mathrm{Only\ VLLM\text{-}based\ Captioning}$ indicate that local parts are described using only vertex-based or VLLM-based captioning, respectively. $\mathrm{w/o\ stage\ 1}$ means that stage 1 is skipped, i.e., no pre-training is conducted for aligning CAD data with textual descriptions. $\mathrm{w/o\ data\ augmentation}$ represents that only the original samples are used during pre-training, without any augmented data. Best performances are in bold.
[…]
Only Vertex-based Captioning 63.6% 1.18 1.02 89.5% 78.3% -
Only VLLM-based Captioning 61.8% 1.26 1.05 89.1% - 64.2%
w/o stage 1 61.3% 1.21 1.16 89.6% 71.5% 60.4%
w/o data augmentation 62.9% 1.18 1.09 88.5% 73.2% 61.8%
Ours 64.9% 1.13 0.98 90.5% 76.4% 65.7%
Effectiveness of the complementary captioning strategy. As shown in Table 2, using only vertex-based or VLLM-based captioning fails to generate complex parts (e.g., a letter V) or simple parts (e.g., a trapezoid), thereby failing to obtain the corresponding Ver-score and VLLM-score. In contrast, the complementary captioning integrating both of them leads to improved performance.
Effectiveness of Pre-training. As depicted in Table 2, omitting stage 1 results in the poorest performance, demonstrating that pre-training is essential for achieving preliminary text-CAD alignment. Additionally, excluding data augmentation during pre-training leads to a performance decline, indicating that diverse augmented samples enhance GeoCAD ’s alignment capability. Together, these findings confirm the effectiveness of the pre-training process.
5 Conclusion
In this paper, we introduce a local geometry-controllable CAD generation method, GeoCAD, enabling users to specify design intent for specific parts through geometric instructions. To the best of our knowledge, GeoCAD is the first to achieve local geometry-control in the CAD generation field. To accomplish this, GeoCAD introduces both vertex-based and VLLM-based captioning pipelines and employs a two-stage training strategy for LLM fine-tuning. Extensive qualitative and quantitative evaluations demonstrate that GeoCAD substantially improves generation quality, validity, and text-to-CAD consistency in local geometry-controllable CAD generation.
Acknowledgement
This work was supported in part by the Key R&D Program of Zhejiang Province (2025C01212), in part by Yongjiang Talent Introduction Programme (2022A-240-G), in part by Ningbo Key R&D Program (2023Z229), in part by The National Nature Science Foundation of China (Grant NOs: 62273303, 62303406, 62273302, 62036009), in part by Ningbo Key R&D Program (NO.: 2025Z055).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of our work in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We fully disclose all the information needed to reproduce the main experimental results of the paper in Sec. 4.1
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code and data are provided in the supplemental material, with sufficient instructions to faithfully reproduce the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Our paper specifies all the training and test details in Sec. 4.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report consistent performance across multiple runs and use fixed random seed settings to support the statistical significance of our results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: For each experiment, the paper provides sufficient information on the computer resources in Sec. 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All existing assets, including code, data, and models, are properly cited in the paper, and their licenses and usage terms are respected in accordance with the original sources.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will release the code and data with accompanying documentation to ensure usability and reproducibility in an anonymous manner.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: For crowdsourcing experiments, the paper includes the full text of instructions given to participants. Workers are paid more than the minimum wage in the country of the data collector.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: There are no potential risks associated with the crowdsourcing experiments.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We describe the usage of LLMs in Sec. 4.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix
Due to space limitations in the main paper, we provide additional results and discussions in this appendix, organized as follows:
  • Sec. A: More Details about VLLM-based Captioning.
  • Sec. B: Detailed Comparison with Existing Work.
  • Sec. C: Detailed Categories of Simple Parts and Complex Parts.
  • Sec. D: Details about Metrics for Evaluating Text-to-CAD consistency.
  • Sec. E: LLMs of Different Scales.
  • Sec. F: Sensitivity Analysis of Key Hyper-parameters in Sampling.
  • Sec. G: Criteria for Dataset Selection.
  • Sec. H: Sketch-level Editing.
  • Sec. I: Failure Cases, Limitations and Future Work.
  • Sec. J: Five-shot Prompt Example.
  • Sec. K: Additional Qualitative Results.
Appendix A More Details about VLLM-based Captioning
The prompt used for VLLM-based captioning is as follows:
"Given a loop in a CAD sketch, provide a brief description of its geometric shape starting with ‘a’ or ‘an’ if identifiable; otherwise, state ‘None’."
Using this prompt, we randomly caption 1k complex local parts with GPT-4o [1] and Qwen2.5-VL-72B-Instruct [4], respectively. Regardless of whether these models output a specific shape or ’None’, we manually evaluate each result by judging its correctness as either "Yes" or "No". The overall captioning accuracy across these 1,000 parts is 91.3% for Qwen2.5-VL-72B-Instruct and 86.5% for GPT-4o. These results indicate that Qwen2.5-VL-72B-Instruct outperforms GPT-4o in this captioning task, which is consistent with with the latest multimodal model leaderboard rankings. Furthermore, given the lower cost of Qwen2.5-VL-72B-Instruct, we use it to caption the remaining complex parts.
Appendix B Detailed Comparison with Existing Work
As mentioned in lines 36-48 in our main paper, existing work struggles to achieve local geometry-controllable CAD generation. Here, we further highlight the differences between CAD-Editor [55], FlexCAD [61] and our GeoCAD. CAD-Editor has difficulty focusing on local generation for two main reasons: 1) It may unintentionally modify the remaining parts, resulting in outputs that do not align with user requirements (as illustrated in the last example of Fig. 1 in the original CAD-Editor paper). 2) CAD-Editor fails to accurately obtain angle and length information, making it incapable of generating even simple parts, such as a right triangle, let alone an isosceles right triangle, as mentioned in line 45 of our main paper. FlexCAD, on the other hand, can focus on local parts but incorporates minimal geometric constraints, thereby struggling to follow geometric instructions. In particular, FlexCAD is unable to understand, let alone follow, simple or complex geometric instructions. This limitation is clearly demonstrated in Fig. 1 of our main paper.
Appendix C Detailed Categories of Simple Parts and Complex Parts
The categories of simple parts include acute triangle, right triangle, obtuse triangle, isosceles triangle, isosceles right triangle (Notably, equilateral triangles do not occur in the DeepCAD [45] dataset), quadrilateral, trapezoid, isosceles trapezoid, kite (Two pairs of adjacent sides equal), parallelogram, rectangle, rhombus, square, circle, semicircle, quarter-circle, three-quarter circle, major-arc loop (defined as containing an arc longer than a semicircle), minor-arc loop (defined as containing an arc shorter than a semicircle), and so on. The remaining local parts are classified as complex, exhibiting more intricate and diverse visual patterns.
Appendix D Details about Metrics for Evaluating Text-to-CAD consistency
As mentioned in lines 213–217 of our main paper, we employ Ver-score, VLLM-score, and Realism to comprehensively evaluate model performance in terms of text-to-CAD consistency. Specifically, to compute Ver-score, we extract vertex coordinates from the generated local parts and analyze their geometric attributes to determine whether they align with the given geometric instructions. To obtain VLLM-score, we first render the local parts into images and then prompt two of the most powerful VLLMs, GPT-4o [1] and Qwen2.5-VL-72B-Instruct [4], to judge whether the rendered images match the corresponding instructions, assigning a binary label: "Yes" or "No." We report the average of their scores in Table 1 of our main paper, where both models significantly outperform the baselines. To evaluate Realism, we randomly render 500 newly generated CAD models into images, with the modified local parts clearly marked. Five crowd workers are then asked to assess whether the generated local parts align with the geometric instructions and do not conflict with the remaining parts. If both criteria are satisfied, they assign a binary label: "Yes"; otherwise, "No." The average score from these workers is reported in Table 1 of our main paper.
Appendix E LLMs of Different Scales
Table 3: Ablation studies on fine-tuning LLMs with different scales. $\mathrm{Llama\text{-}3\text{-}8B}$ is the model used in our main paper to enable a fair comparison with FlexCAD [61]. $\mathrm{Transformer\text{-}4M}$ is a small Transformer-based [42] language model, with a total number of trainable parameters comparable to that of our model in the main paper when using LoRA. $\mathrm{Llama\text{-}3\text{-}8B\text{-}Full}$ denotes full-parameter fine-tuning. $\mathrm{Llama\text{-}3\text{-}8B}$, $\mathrm{Qwen2.5\text{-}3B\text{-}Instruct}$, and $\mathrm{Qwen2.5\text{-}7B\text{-}Instruct}$ are all fine-tuned using LoRA. The best results are shown in bold, and the second-best results are marked with *.
[…]
Transformer-4M 59.1% 1.32 1.26 85.5% 69.3% 51.2%
Llama-3-8B-Full 67.5%* 1.02* 1.06 89.7% 78.9%* 64.2%
Llama-3-8B 64.9% 1.13 0.98* 90.5% 76.4% 65.7%*
Qwen2.5-3B-Instruct 65.8% 1.01 1.10 87.4% 74.2% 64.9%
Qwen2.5-7B-Instruct 68.7% 1.05 0.86 90.1%* 79.8% 70.2%
As shown in Table 3, $\mathrm{Transformer\text{-}4M}$ achieves the lowest performance, confirming that LLMs play a key role in enhancing local CAD generation. $\mathrm{Llama\text{-}3\text{-}8B\text{-}Full}$ performs comparably to $\mathrm{Llama\text{-}3\text{-}8B}$, demonstrating the effectiveness of the LoRA strategy [16]. As two of the most popular open-source LLMs, $\mathrm{Qwen2.5\text{-}7B\text{-}Instruct}$ slightly outperforms $\mathrm{Llama\text{-}3\text{-}8B}$.
Appendix F Sensitivity Analysis of Key Hyper-parameters in Sampling
Table 4: Effectiveness analysis of key hyper-parameters, including the sampling temperature $\tau$ and $\mathrm{Top\text{-}p}$. Best performances are in bold and the second-bests are marked by *.
[…]
$\tau=0.7$ 63.4% 1.18 1.03 91.2% 75.9% 63.2%
$\tau=0.9$ 64.9%* 1.13 0.98* 90.5%* 76.4%* 65.7%
$\tau=1.1$ 65.6% 1.16* 0.95 89.1% 77.5% 65.1%*
$\mathrm{Top\text{-}p}=0.8$ 64.1% 1.21 1.09 91.0% 75.3% 64.4%
$\mathrm{Top\text{-}p}=0.9$ 64.9%* 1.13 0.98* 90.5%* 76.4%* 65.7%*
$\mathrm{Top\text{-}p}=1.0$ 65.2% 1.18* 0.92 88.3% 76.9% 66.8%
As shown in Table 4, we conduct a sensitivity analysis on key hyperparameters, including the sampling temperature $\tau$ and $\mathrm{Top\text{-}p}$. All other experimental settings follow those described in Section 4.2 of our main paper. In general, increasing $\tau$ or $\mathrm{Top\text{-}p}$ results in more diverse and stochastic predictions. However, this comes at the cost of reduced $\mathrm{PV}$, while other metrics tend to improve, consistent with findings in [61]. In our experiments, we balance this trade-off by selecting $\tau$ and $\mathrm{Top\text{-}p}$ values that ensure the $\mathrm{PV}$ remains above 90%.
Appendix G Criteria for Dataset Selection
DeepCAD [45] is a suitable dataset for evaluation, and the reasons are detailed below: 1) Scale: DeepCAD is a large-scale 3D CAD dataset, comprising over 178k samples. 2) Relevance to Controllability: Compared to 2D sketch datasets, DeepCAD better reflects the requirements of controllable generation, as aligning local parts within 3D CAD models is both more challenging and more practical. 3) Design Process Alignment: In contrast to other 3D CAD datasets, such as the ABC dataset [19], DeepCAD includes sketch-and-extrusion sequences that closely mirror the design workflows of commercial CAD tools like SolidWorks and AutoCAD. 4) Community Adoption: Due to its characteristics, DeepCAD is also the only choice for prior studies, including SkexGen [50], HNC-CAD [48], CAD-Editor [55], CADFusion [44], Text2CAD [18], and FlexCAD [61].
Figure A1: An example of sketch-level editing.
Appendix H Sketch-level Editing
For sketch-level editing, if a sketch contains multiple loops, ideally, we would like to learn the inter-loop constraints (e.g., symmetry, patterns, etc.) that define the overall structure. However, as mentioned above, DeepCAD is currently the only dataset suitable for controllable 3D CAD generation, and unfortunately, such inter-loop constraint annotations are not provided in the dataset. Fortunately, even without supervision from these constraints, sketch-level editing is still achievable based on our loop-level editing capability. This is because the loop serves as the fundamental element of a sketch. For example, as shown in Fig. A1, if a user selects a sketch consisting of two symmetric loops and wishes to replace them with another pair of symmetric loops, the following automatic process can be performed: 1) Estimate the center point of each original loop by averaging its coordinate points, which are extracted using string matching. 2) Determine the symmetry axis based on the two center points. 3) Generate a new local loop through GeoCAD replacing one of the orignal loops. 4) Reflect the newly generated loop across the symmetry axis to produce the second symmetric loop, thereby replacing both original loops.
Appendix I Failure Cases, Limitations and Future Work
Figure A2: Failure cases. The generated local parts align well with the user’s geometric instructions but do not integrate smoothly with the remaining parts of the original CAD model.
Failure cases. Despite notable advancements, our GeoCAD sometimes results in failure cases. As shown in Fig. A2, given a CAD model, when only the special part is modified (i.e., the part upon which the remaining parts are constructed and strictly aligned in size), the unchanged remaining parts may lead to structural conflicts with it. To mitigate this issue, when modifying the special parts, users should provide geometric instructions that account for the constraints imposed by the remaining parts, since the DeepCAD dataset does not annotate the relationships between different parts.
Limitations and future work. In this paper, we fine-tune LLMs to enable local geometry-controllable CAD generation, primarily guided by textual instructions. However, in practice, certain complex local parts may be difficult or even impossible to describe using text alone. Thus, in the future, if users can complement textual inputs with hand-drawn images for local geometry-controllable CAD generation, they may be able to convey their design intent more effectively. Given the strong capabilities of VLLMs in both CAD generation and text understanding, our future work aims to develop a more advanced multimodal LLM tailored for controllable CAD generation from both text and image inputs.
Appendix J Five-shot Prompt Example
To better illustrate the implementation details of the baselines and our GeoCAD in Table 1 of our main paper, we present a five-shot prompt example, as shown in Fig. A3.
Figure A3: A five-shot prompt example used in Table 1 of our main paper.
Figure A4: Additional Qualitative Results.
Appendix K Additional Qualitative Results
We provide additional qualitative results in Fig. A4.